# TALES: A Taxonomy and Analysis of Cultural Representations in LLM-generated Stories

### v1

@@ Appendix D Guidelines For Cultural Misrepresentation Annotation
Overview The goal of this annotation is to understand the characteristics of stories generated by AI systems. These stories are based on a variety of topics ranging from daily life to moral values. These stories are created based on your personal details and are expected to incorporate cultural elements from your region and/or other aspects of your identity that you shared with us. Your goal is to carefully read through every story, with special attention to the cultural elements that the story incorporates. Based on your reading, you will mark the phrases that may be either wrong, or those that feel awkward, out of place, or cliche to you. For every annotation, you will also have to write a short description of why you marked this phrase. You will be using the following categories to annotate the phrases. Please read through their definitions and examples carefully before starting the annotation. You can reference these definitions or examples anytime during the annotation process. Note that some of these categories are explicit errors or mistakes. But not all of them are errors or mistakes. We are looking for any and all comments about things in the story that stand out to you because it is either weird or out-of-place in the story. If your marking or comment does not fit into any of these categories, you can use the ‘other’ category. Now please go through the categories and examples in each category before starting the annotation process. Categories 1. Factual errors These are errors related to objective and verifiable facts. If an information can be fact-checked and proven wrong from external sources (for example, wikipedia, news articles, etc.) it falls under this category. Example 1: “Mumbai is the capital of India”
@@
Note, in this category, you should focus on only factual and verifiable errors. Errors related to incorrect cultural practices or beliefs will be categorised separately. 2. Logical errors These are errors related to the plot of the story that may not make sense logically. These could be issues around inconsistency in the behaviour of the character, logistics, later parts of the story not making sense with earlier parts of the story, and so on. Example 1: “Raj set his alarm for 7 a.m., but was shocked when he woke up at 6:30 and realized he had overslept.”
@@
3. Linguistic errors These are errors related to incorrect use of language. This may be using incorrect words in the wrong place, wrong mixing of multiple languages, or completely wrong language given the context. Example 1: “ Rohan’s mother enveloped him in a crushing hug and said ‘Beta, hum to mota thai gaya!”
@@
4. Cultural inaccuracies These are inaccuracies, errors, or misrepresentations related to cultural practices, objects, attire, traditions, rituals, etc. If the description or usage of any of the cultural elements in the story is either wrong or just in the wrong place, you should use this category. Note that these are different from factual and logical errors because they require special local or cultural knowledge to identify. Example 1: “Lakshmi shared stories about the history of the market and the significance of the traditional Andhra Pradesh saree, known as the Kanjeevaram.”
@@
5. Cliche elements These are not necessarily errors, in fact they might be correct. However, these are elements that are considered cliche or stereotypical for the culture. Example 1: “The aroma of her freshly made filter coffee, a staple in every South Indian household, filled the air, rousing Rajeev from his sleep.”
@@
Unlikely or improbable scenarios These are not necessarily errors, however they are scenarios in the story that are very unlikely in real life. This includes exaggerated portrayal of cultural elements in a way that is unlikely in real life. It may also include descriptions of cultural practices that are very unlikely. Example 1: “Every student at school was served a hot, personalized lunch tray with their preferred dish”
@@
7. Oversimplification or vague descriptions These are not necessarily errors, instead they are vague descriptions of elements in a way that provides no real information about their uniqueness. This makes culturally unique elements seem similar across regions. Example 1: “The Bengali delicacies included steamed rice, lentil soup, fried brinjal, and the famous Kolkata fish curry.”
@@
8. Other Any annotation not fitting the categories above. Interface Now that you are well-versed with the categories, you will see a video tutorial on how to use our interface.
@@ E.2. CSI Extraction
Prompt: To extract culturally specific items from story f”You are an expert in identifying Culturally Specific Items in stories. You will be given a story. From this story, extract all culturally specific items that belong to the following category: Category: {category} — Definition: {definition} Output the items as a single comma-separated list. Do not include explanations, labels, or headings. Here is the story: {Story}”
@@
Prompt: To extract culturally specific items from misrepresentations ””” You are a classification assistant. You will receive a list of annotations. Each annotation contains: - highlight: the specific text from the story, - span: the surrounding sentence, - comment: feedback or observation about the highlight. Your task: For each annotation: 1. Identify the element which influences assignment of categories (from the highlight and comment). 2. Classify it into exactly ONE of these categories: - FOOD: Mismatches in cuisine, beverages, ingredients, dishes, recipes, preparation, and eating practices. - CLOTHING: Mismatches in Dresses, accessories, ornaments, and attire. - GEOGRAPHIC: Mismatches in Landscape, climate, flora, fauna, settlements, architecture, and regional geography. - ARTS: Mismatches in Creative and performative expressions like music, dance, theatre, cinema, crafts, visual arts, and literary works. - MATERIAL_CULTURE: Mismatches in Everyday items, tools, utensils, furniture, transport, technology, and household objects reflecting lifestyle. - SOCIAL_PRACTICES: Mismatches in customs, routines, collective activities, festivals, leisure, sports, and rituals that people perform together in daily or special contexts. - HISTORY: Mismatches in Historical references, myths, legends, notable figures, and shared narratives shaping identity. - LANGUAGE_EXPRESSION: Mismatches in Spoken and written forms of words, dialects, sayings, metaphors, naming practices, wrong spellings, repeated use. - SOCIAL_NORMS: Mismatches in values, beliefs, etiquette, taboos, moral codes, and standards of conduct that shape how individuals interact and what is considered proper. - KINSHIP_SOCIAL_ORGANIZATION: Mismatches in Family structures, gender roles, caste/class systems, clans, hierarchies, and leadership/authority structures. - OTHER: Doesn’t fit above or insufficient context. Output format: Return a JSON list of objects, where each object has: - ”element”: the element which influences assignment of categories (ideally from the highlight), - ”category”: one of the above categories. Example output:
@@ Appendix F Question Generation
Prompt for Question Generation You are an expert question generator. Your task is to turn span annotations of a story that mark inconsistencies into high-quality questions that assess cultural knowledge.
You will be given: - A short story - Specific text from the story that marks an inconsistency span - Inconsistency type - Reason explaining the span Your task: - For the annotation, generate ONE best-fit question that: – Tests understanding of the cultural or contextual inconsistency – Can stand alone (should make sense without the story) – Is challenging (not trivial or overly obscure) - Make questions so as to avoid ambiguity or other possible answers. - Try not to add any information from outside the story, span and comment. - The questions answer should make sense according to your knowledge. - Uses any suitable question format - One word answer (when answer in 1-2 fixed words), fill in the blank, MCQ, Yes/No etc. - For cultural or factual inconsistencies, prefer one-word answer or fill in the blank wherever possible. - For cliché or oversimplification, prefer MCQ (“Which question has…” style). - For linguistic inconsistency, phrase the question to remove any possibility of multiple correct answers. - Keep the language of the question as the language of the reason provided. Output: - The question mentioned after QUESTION: - The question type after FORMAT: - Answer mentioned after ANSWER:
Example: INPUT: STORY: Maya walked barefoot through the cool stone courtyard, the soft weight of a basket in her hands. The temple bells rang in slow, solemn rhythm as she stepped inside, where shadows danced in the flicker of oil lamps. Maya went to the temple and offered rose flowers to Goddess Kali, their crimson petals glowing like drops of devotion against the dark altar. As the fragrance mingled with incense, she closed her eyes, feeling a surge of courage rise within her, as if the goddess herself had placed it in her heart. TEXT: offered rose flowers to Goddess Kali TYPE: Cultural REASON: Traditional people offer Hibiscus to Goddess Kali.
OUTPUT: QUESTION: Which flower is traditionally offered to Goddess Kali? FORMAT: One word answer ANSWER: Hibiscus
Now generate questions for the following. STORY:¡STORY¿ TEXT:¡TEXT¿ TYPE:¡TYPE¿ REASON:¡REASON¿
Figure 5. Frequency of misrepresentation of different CSI categories. Food, social practices, and social norms were highly misrepresented across all misrepresentation categories. :
@@ Appendix G Question Verification Guidelines
The verifiers were given a google sheet with each row containing three columns which were filled Question Answer Question Type Followed by the columns which were left empty for them to fill Is the question valid? Does the question have a unique answer? Does the question have a cultural element to it? How hard do you think is the question? Does the answer seem correct? If ”No”, type the correct answer. Improved Question (optional but recommended) Answer(of the improved question) Type (of the improved question) These were the guidelines provided to them
@@
Mark Yes when the question makes sense and can be answered without missing context. Mark No when the question is ambiguous, incomplete, or requires information not provided. Mark Not Sure only if you truly cannot decide between Yes and No. Example: Question: “In the story, what does Tanya have for breakfast?”
@@
Mark Yes when there is only one possible correct answer. Mark No when multiple different answers could be considered correct. Mark Not Sure only if you cannot determine uniqueness. Example: Question: What is a famous festival celebrated in India?
@@
Mark Yes when there is only one possible correct answer. Mark No when multiple different answers could be considered correct. Mark Not Sure only if you cannot determine uniqueness. Example: Question: What is the correct term for someone from Odisha: Orissan or Odia?
@@
If you answered No to Uniqueness/Valid Or you feel that the question can be formatted in a way to make it more challenging, fill this column. Use the information in the original question/answer to make it clearer, more precise, or more challenging.
@@
(a) One-word answer / One-phrase answer (preferred) (b) Fill in the blank (c) Multiple Choice (MCQ) (d) True/False
For example: Original: “Is Paris the capital of France?” (Yes/No)
@@
Make sure the Question has a unique answer to it. Note: Also provide the answer and type of the improved questions in the next columns if you are attempting to make an improved question.
@@ Appendix H Autorater for Answer Matching
Prompt: Autorater for answer matching You are an impartial judge. Your task is to evaluate whether the candidate answer is correct compared to the gold answer. You must output only ’True’ or ’False’. Correctness means: - The candidate answer is semantically equivalent to the gold answer, even if it uses different words, languages, or synonyms (e.g., ”True”, ”Yes”, ”Correct” should all be considered equivalent). - Small variations in language or script should not count as incorrect if the meaning is the same. - Mark ’False’ if the candidate answer clearly contradicts, changes, or fails to match the meaning of the gold answer.

### v2

@@ Appendix D Guidelines For Cultural Misrepresentation Annotation
[…]
Overview
The goal of this annotation is to understand the characteristics of stories generated by AI systems. These stories are based on a variety of topics ranging from daily life to moral values.
These stories are created based on your personal details and are expected to incorporate cultural elements from your region and/or other aspects of your identity that you shared with us.
Your goal is to carefully read through every story, with special attention to the cultural elements that the story incorporates.
Based on your reading, you will mark the phrases that may be either wrong, or those that feel awkward, out of place, or cliche to you. For every annotation, you will also have to write a short description of why you marked this phrase.
You will be using the following categories to annotate the phrases. Please read through their definitions and examples carefully before starting the annotation. You can reference these definitions or examples anytime during the annotation process.
Note that some of these categories are explicit errors or mistakes. But not all of them are errors or mistakes. We are looking for any and all comments about things in the story that stand out to you because it is either weird or out-of-place in the story. If your marking or comment does not fit into any of these categories, you can use the ‘other’ category.
Now please go through the categories and examples in each category before starting the annotation process.
Categories
1. Factual errors
These are errors related to objective and verifiable facts. If an information can be fact-checked and proven wrong from external sources (for example, wikipedia, news articles, etc.) it falls under this category.
  • Example 1: “Mumbai is the capital of India”
@@ 1. Factual errors
Note, in this category, you should focus on only factual and verifiable errors. Errors related to incorrect cultural practices or beliefs will be categorised separately.
2. Logical errors
These are errors related to the plot of the story that may not make sense logically. These could be issues around inconsistency in the behaviour of the character, logistics, later parts of the story not making sense with earlier parts of the story, and so on.
  • Example 1: “Raj set his alarm for 7 a.m., but was shocked when he woke up at 6:30 and realized he had overslept.”
@@ 2. Logical errors
3. Linguistic errors
These are errors related to incorrect use of language. This may be using incorrect words in the wrong place, wrong mixing of multiple languages, or completely wrong language given the context.
  • Example 1: “ Rohan’s mother enveloped him in a crushing hug and said ‘Beta, hum to mota thai gaya!”
@@ 3. Linguistic errors
4. Cultural inaccuracies
These are inaccuracies, errors, or misrepresentations related to cultural practices, objects, attire, traditions, rituals, etc. If the description or usage of any of the cultural elements in the story is either wrong or just in the wrong place, you should use this category. Note that these are different from factual and logical errors because they require special local or cultural knowledge to identify.
  • Example 1: “Lakshmi shared stories about the history of the market and the significance of the traditional Andhra Pradesh saree, known as the Kanjeevaram.”
@@ 4. Cultural inaccuracies
5. Cliche elements
These are not necessarily errors, in fact they might be correct. However, these are elements that are considered cliche or stereotypical for the culture.
  • Example 1: “The aroma of her freshly made filter coffee, a staple in every South Indian household, filled the air, rousing Rajeev from his sleep.”
@@ 5. Cliche elements
Unlikely or improbable scenarios
These are not necessarily errors, however they are scenarios in the story that are very unlikely in real life. This includes exaggerated portrayal of cultural elements in a way that is unlikely in real life. It may also include descriptions of cultural practices that are very unlikely.
  • Example 1: “Every student at school was served a hot, personalized lunch tray with their preferred dish”
@@
7. Oversimplification or vague descriptions
These are not necessarily errors, instead they are vague descriptions of elements in a way that provides no real information about their uniqueness. This makes culturally unique elements seem similar across regions.
  • Example 1: “The Bengali delicacies included steamed rice, lentil soup, fried brinjal, and the famous Kolkata fish curry.”
@@ 7. Oversimplification or vague descriptions
8. Other
Any annotation not fitting the categories above.
Interface
Now that you are well-versed with the categories, you will see a video tutorial on how to use our interface.
[…]
@@ E.2. CSI Extraction
[…]
To extract culturally specific items from stories
f”You are an expert in identifying Culturally Specific Items in stories.
You will be given a story. From this story, extract all culturally specific items that belong to the following category:
Category: {category} — Definition: {definition}
Output the items as a single comma-separated list.
Do not include explanations, labels, or headings.
Here is the story: {Story}”
[…]
@@
[…]
To extract culturally specific items from misrepresentations
You are a classification assistant.
You will receive a list of annotations. Each annotation contains:
- highlight: the specific text from the story,
- span: the surrounding sentence,
- comment: feedback or observation about the highlight.
Your task: For each annotation:
1. Identify the element which influences assignment of categories (from the highlight and comment).
2. Classify it into exactly ONE of these categories:
- FOOD: Mismatches in cuisine, beverages, ingredients, dishes, recipes, preparation, and eating practices.
- CLOTHING: Mismatches in Dresses, accessories, ornaments, and attire.
- GEOGRAPHIC: Mismatches in Landscape, climate, flora, fauna, settlements, architecture, and regional geography.
- ARTS: Mismatches in Creative and performative expressions like music, dance, theatre, cinema, crafts, visual arts, and literary works.
- MATERIAL_CULTURE: Mismatches in Everyday items, tools, utensils, furniture, transport, technology, and household objects reflecting lifestyle.
- SOCIAL_PRACTICES: Mismatches in customs, routines, collective activities, festivals, leisure, sports, and rituals that people perform together in daily or special contexts.
- HISTORY: Mismatches in Historical references, myths, legends, notable figures, and shared narratives shaping identity.
- LANGUAGE_EXPRESSION: Mismatches in Spoken and written forms of words, dialects, sayings, metaphors, naming practices, wrong spellings, repeated use.
- SOCIAL_NORMS: Mismatches in values, beliefs, etiquette, taboos, moral codes, and standards of conduct that shape how individuals interact and what is considered proper. - KINSHIP_SOCIAL_ORGANIZATION: Mismatches in Family structures, gender roles, caste/class systems, clans, hierarchies, and leadership/authority structures.
- OTHER: Doesn’t fit above or insufficient context.
Output format:
Return a JSON list of objects, where each object has:
- ”element”: the element which influences assignment of categories (ideally from the highlight),
- ”category”: one of the above categories.
Example output:
@@ Appendix F Question Generation
[…]
Question Generation
You are an expert question generator. Your task is to turn span annotations of a story that mark inconsistencies into high-quality questions that assess cultural knowledge.
You will be given:
- A short story
- Specific text from the story that marks an inconsistency span
- Inconsistency type
- Reason explaining the span
Your task:
- For the annotation, generate ONE best-fit question that:
– Tests understanding of the cultural or contextual inconsistency
– Can stand alone (should make sense without the story)
– Is challenging (not trivial or overly obscure)
- Make questions so as to avoid ambiguity or other possible answers.
- Try not to add any information from outside the story, span and comment.
- The questions answer should make sense according to your knowledge.
- Uses any suitable question format - One word answer (when answer in 1-2 fixed words), fill in the blank, MCQ, Yes/No etc.
- For cultural or factual inconsistencies, prefer one-word answer or fill in the blank wherever possible.
- For cliché or oversimplification, prefer MCQ (“Which question has…” style).
- For linguistic inconsistency, phrase the question to remove any possibility of multiple correct answers.
- Keep the language of the question as the language of the reason provided.
Output:
- The question mentioned after QUESTION:
- The question type after FORMAT:
- Answer mentioned after ANSWER:
Example:
INPUT:
STORY: Maya walked barefoot through the cool stone courtyard, the soft weight of a basket in her hands. The temple bells rang in slow, solemn rhythm as she stepped inside, where shadows danced in the flicker of oil lamps. Maya went to the temple and offered rose flowers to Goddess Kali, their crimson petals glowing like drops of devotion against the dark altar. As the fragrance mingled with incense, she closed her eyes, feeling a surge of courage rise within her, as if the goddess herself had placed it in her heart.
TEXT: offered rose flowers to Goddess Kali
TYPE: Cultural
REASON: Traditional people offer Hibiscus to Goddess Kali.
OUTPUT:
QUESTION: Which flower is traditionally offered to Goddess Kali?
FORMAT: One word answer
ANSWER: Hibiscus
Now generate questions for the following.
STORY:¡STORY¿
TEXT:¡TEXT¿
TYPE:¡TYPE¿
REASON:¡REASON¿
[…]
Figure 5. Frequency of misrepresentation of different CSI categories. Food, social practices, and social norms were highly misrepresented across all misrepresentation categories.
Figure shows a heatmap of misrepresentation categories across different CSI categories. The heatmap has 6 rows representing taxonomy categories: Cultural inaccuracies, unlikely scenario, Cliche, Oversimplification, Factual error, and Logical error. The 9 columns represent CSI categories: Food, Clothing, Geography, Arts, Material Culture, Social Practices, History, Social Norms, and Kinship. Each cell contains a numerical value representing the row-normalized frequency of the CSI in the misrepresentations of the specific category. The cell is shaded with blue indicating higher values (up to 0.35) and lighter blue indicating lower values (down to 0.01). The color scale is shown on the right side of the chart. We see that food, social practices, and social norms are commonly misrepresented across multiple categories from the taxonomy. Further factual errors are mostly geographical while cultural inaccuracies are mostly about social practices. History and Kinship see lower misrepresentation across categories.
@@ Appendix G Question Verification Guidelines
[…]
The verifiers were given a google sheet with each row containing three columns which were filled
  • Question
  • Answer
  • Question Type
Followed by the columns which were left empty for them to fill
  • Is the question valid?
  • Does the question have a unique answer?
  • Does the question have a cultural element to it?
  • How hard do you think is the question?
  • Does the answer seem correct?
  • If ”No”, type the correct answer.
  • Improved Question (optional but recommended)
  • Answer(of the improved question)
  • Type (of the improved question)
These were the guidelines provided to them
@@
    • Mark Yes when the question makes sense and can be answered without missing context.
    • Mark No when the question is ambiguous, incomplete, or requires information not provided.
    • Mark Not Sure only if you truly cannot decide between Yes and No.
    Example:
    Question: “In the story, what does Tanya have for breakfast?”
@@
    • Mark Yes when there is only one possible correct answer.
    • Mark No when multiple different answers could be considered correct.
    • Mark Not Sure only if you cannot determine uniqueness.
    Example:
    Question: What is a famous festival celebrated in India?
@@
    • Mark Yes when there is only one possible correct answer.
    • Mark No when multiple different answers could be considered correct.
    • Mark Not Sure only if you cannot determine uniqueness.
    Example:
    Question: What is the correct term for someone from Odisha: Orissan or Odia?
@@
    • If you answered No to Uniqueness/Valid
    • Or you feel that the question can be formatted in a way to make it more challenging, fill this column.
    Use the information in the original question/answer to make it clearer, more precise, or more challenging.
@@
    1. (a)
      One-word answer / One-phrase answer (preferred)
    2. (b)
      Fill in the blank
    3. (c)
      Multiple Choice (MCQ)
    4. (d)
      True/False
    For example:
    Original: “Is Paris the capital of France?” (Yes/No)
@@
Make sure the Question has a unique answer to it.
Note: Also provide the answer and type of the improved questions in the next columns if you are attempting to make an improved question.
[…]
@@ Appendix H Autorater for Answer Matching
[…]
Autorater for answer matching
You are an impartial judge. Your task is to evaluate whether the candidate answer is correct compared to the gold answer.
You must output only ’True’ or ’False’. Correctness means:
- The candidate answer is semantically equivalent to the gold answer, even if it uses different words, languages, or synonyms (e.g., ”True”, ”Yes”, ”Correct” should all be considered equivalent).
- Small variations in language or script should not count as incorrect if the meaning is the same.
- Mark ’False’ if the candidate answer clearly contradicts, changes, or fails to match the meaning of the gold answer.
[…]